Add and simplify
x/(a+b) + y(b+a)
Add
(2/x) + (3/-x)
How do I find the numerator of the sum
((y+3)/y) + (y/(y+3))
Add and Simplify
9/(3-w) +w^2/(w-3)
Add (3/x)+(4/y)
Subtract (5/x)- ((x+3))/x^2)

Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

# 1



# 2




# 3



# 4




# 5

Answer by stanbon(75887)

Add and simplify
x/(a+b) + y(b+a)
--
The denominators are the same so add the numberators to get:
= (x+y)/(a+b)
-----------------------------
Add
(2/x) + (3/-x)
---
Modify:
(2/x) - (3/x)
= (2-3)/x
= -1/x
-----------------------------
How do I find the numerator of the sum
((y+3)/y) + (y/(y+3))
Least common denominator: y(y+3)
= [(y+3)^2 + y^2] /[y(y+3)]
= [2y^2 + 6y + 4] / [y(y+3)]
= [2(y+2)(y+1)] / [y(y+3)]
-------------------------------------------------
Add and Simplify
9/(3-w) + w^2/(w-3)
---
Modify:
-9/(w-3) + w^2/(w-3)
---
Combine the numerators over the common denominator:
= (w^2-9)/(w-3)
---
= (w+3)
--------------------------------------------------
Add (3/x)+(4/y)
---
least common denominator: xy
--
Modify:
(3y/xy) + (4x/xy
Combine:
(3y+3x)/xy
--------------------------------------------------
Subtract (5/x)- ((x+3))/x^2)
--
least common denominator: x^2
---
Modify:
(5x/x^2) + ((x+3)/x^2
---
Combine:
[5x + x + 3]/(x^2)
----
Simplify:
(6x+3)/x^2
====================================================
Cheers,
Stan H.